# On the Zero-Error Capacity of Semantic Channels with Input and Output Memories

### v1

@@ II-A Channel with memory
    The channel inputs have been the same (e.g., 0) for the last $K_{1}$ successive epochs, but a different input (e.g., $x_{t}=1$) is observed in the current slot.
@@
    The channel outputs have been the same (e.g., 0) for the last $K_{2}$ successive epochs, but a different input (e.g., $x_{t}=1$) is observed in the current slot.
To illustrate the motivation behind this channel, consider an analogy from semantic communication [10, 11]. Imagine a conversation between two parties, Alice and Bob. Alice says $K_{1}$ sentences about the weather and then suddenly says a sentence about emotions. Bob hesitates, unsure if the last sentence is still about the weather. Conversely, if Bob interprets the last $K_{2}$ sentences said by Alice as being about the weather but suddenly hears a sentence about emotions, he also hesitates, unsure whether the last sentence is still about the weather or not.
@@ II-B Zero-error capacity
Note that the base of the logarithm in (7) is $2$, and will be omitted throughout this paper. Since $0\leq|\mathcal{B}^{*}_{n}|\leq|\mathcal{X}|^{n}=2^{n}$, we have $0\leq C(M_{K_{1},K_{2}})\leq 1$.
@@ III-C The $K_{1},K_{2}\geq 2$ case
An illustration of the bounds in (28) is given in Fig. 2. As can be seen, with the increase in $K_{1}$, the upper and lower bounds of $C(M_{K_{1},K_{2}})$ gradually converge.

### v2

@@ II-A Channel with memory
    The channel inputs have been the same (e.g., 0) for the last $K_{1}-1$ successive epochs, but a different input (e.g., $x_{t}=1$) is observed in the current slot.
@@
    The channel outputs have been the same (e.g., 0) for the last $K_{2}-1$ successive epochs, but a different input (e.g., $x_{t}=1$) is observed in the current slot.
To illustrate the motivation behind this channel, consider an analogy from semantic communication [10, 11] inspired by applications like event detection in sensor networks[12] and anomaly detection in security systems[13]. Imagine two parties, Alice and Bob: Alice provides a series of reports (inputs), and Bob interprets each one (outputs). Alice might report stable environmental readings or normal system status, and Bob interprets these as a stable state. Suddenly, Alice reports a significant deviation, such as a temperature spike or unusual movement. Bob must decide if this is a meaningful anomaly or a minor fluctuation. To make this decision, he considers both Alice’s recent reports and his own prior interpretations. The semantic channel’s structure captures how past inputs (Alice’s reports) and past outputs (Bob’s interpretations) influence the current transmission. Zero-error transmission via this channel helps ensure that Bob’s interpretations remain accurate, providing a reliable response.
@@ II-B Zero-error capacity
Note that the base of the logarithm in (7) is $2$, and will be omitted throughout this paper. Since $1\leq|\mathcal{B}^{*}_{n}|\leq|\mathcal{X}|^{n}=2^{n}$, we have $0\leq C(M_{K_{1},K_{2}})\leq 1$.
In the example with Alice and Bob, each report Alice sends represents a codeword in the channel, and all possible reports form a zero-error code. The zero-error property ensures that each codeword transmitted has a unique interpretation.
@@ III-C The $K_{1},K_{2}\geq 2$ case
$K_{2}$apacity369121500.51
Figure 2: Zero-error capacity of the semantic channel when $K_{1}=1$ or $K_{2}=2,3$ or $K_{1}\geq K_{2}\geq 4$.
$K_{1}$apacity369121500.51upper boundlower bound
@@
[…]
Figure 4: The hollow points represent the cases where the zero-error capacities are determined. The solid points represent the cases where the zero-error capacities are bounded.
@@
Fig. 2 presents the zero-error capacity of the semantic channel for cases where $K_{1}=1$ or $K_{2}=2,3$ or $K_{1}\geq K_{2}\geq 4$. An illustration of the bounds in (27) and (28) is given in Fig. 3. As can be seen, with the increase in $K_{1}$, the upper and lower bounds of $C(M_{K_{1},K_{2}})$ gradually converge. Additionally, Fig. 4 provides a comprehensive illustration of both determined zero-error capacities, and undetermined but bounded zero-error capacities.